# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@ B.3 Algorithm of Data Structure
Lemma B.3 (Runtime of initialization, Algorithm 1).
@@
Lemma B.4 (Runtime of query, Algorithm 2).
@@
Lemma B.5 (Privacy of query, Algorithm 2).
The output of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
@@
Suppose that our dataset is $X\in[-R,R]^{n}$. Note that we only add noise into the pre-processing stage. There is no noise in the query stage. Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma 6.3). Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension. We will then blow up the DP parameter by $\log n$ factor. Thus, using the basic composition rule (Fact 3.10), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP. ∎
@@
Lemma B.6 (Accuracy of query, Algorithm 2).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 3.7, we have the error of Query(see Algorithm 2) output is upper bounded by:
@@
We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$. From Fact 3.7, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
@@
Using Lemma A.3, we can have
@@ B.4 Disjoint Intervals
Lemma B.7 (Sum of disjoint interval queries).
@@
Then, we have $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
@@
From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output. Then, from Fact 3.11 and the disjoint intervals in Algorithm 7, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2 for interval $S_{j}$. Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$. From the proof of Lemma B.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
@@
Using Lemma A.3, we can have error bounded by
@@
Lemma B.8 (Weighted sum of disjoint interval queries, formal version of Lemma 6.2).
@@
Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
@@
The DP proof is the same as in the proof of Lemma B.7.
Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma B.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$. From the proof of Lemma B.7, we know the variance of $z_{j}$ is bounded by
@@
Using Lemma A.3, we can have error bounded by
@@
Appendix C Weighted $\ell_{p}^{p}$ Distance
In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea. In Section C.1, we show how to solve one dimensional weighted problem. In Section C.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
@@ Appendix C Weighted $\ell_{p}^{p}$ Distance
C.1 One Dimensional Weighted Distance
@@
Lemma C.1 (Weighted distance one dimension).
@@
C.2 High Dimensional Weighted Distance
@@
Lemma C.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma 6.4).
@@
where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma C.1.
Appendix D One-Dimensional Weighted $\ell_{1}$ Distance Query
In this section, we generalize the algorithms in [8] to weighted distance. Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e. $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$. In Section D.1, we analyze the runtime of our algorithm. In Section D.2, we analyze the DP and accuracy of our algorithm. In Section D.3, we give the theorem for our DPTreeDistance data structure.
@@ Appendix D One-Dimensional Weighted $\ell_{1}$ Distance Query
1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
@@
3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 1
@@
7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma C.1
@@
13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 1, Lemma 6.3
@@
1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma D.2, Lemma D.4, Lemma D.5
3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 6
@@ D.1 Runtime Analysis
Lemma D.1 (Runtime of initialization, Algorithm 6).
@@
Lemma D.2 (Runtime of DistanceQuery, Algorithm 7).
@@
Remark D.3.
In Line 8 and 13 of Algorithm 7, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma C.1, i.e. $|k-j|R/n$. This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
D.2 Privacy and Accuracy Analysis
@@
Lemma D.4 (Privacy of DistanceQuery, Algorithm 7).
The output of DistanceQuery (Algorithm 7) is $(\epsilon,\delta)$-DP.
@@
From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output. We observe that intervals in Algorithm 7 are disjoint. Then, following the same logic in the proof of Lemma B.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP. ∎
@@
Lemma D.5 (Accuracy of DistanceQuery, Algorithm 7).
@@
From Lemma B.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
@@
Then from the proof of Lemma B.8, we can know that the variance is given by
@@
since the sensitivity $\Delta=2R_{w}$ from Lemma 6.3.
Using Lemma A.3, we can have additive error bounded by
@@
D.3 One Dimension Single Data Structure
@@
Theorem D.6 (DPTreeDistance data structure, formal version of Theorem 6.6).
There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
@@
Appendix E High-Dimensional Weighted $\ell_{1}$ Query
In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from [8]. In Section E.1, we give the analysis of Algorithm 8. In Section E.2, we give the theorem of our DPTreeHighDim data structure.
Algorithm 6,7 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function. We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
@@
1:datastrucutre DPTreeHighDim $\triangleright$ Theorem E.3
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 6
@@
7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma C.2
@@
11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 6
@@
14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma E.1, Lemma E.2
@@
17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 7
@@
Lemma E.1 (Privacy of DistanceQuery, Algorithm 8).
@@
  • Let $\epsilon\leq O(\log(1/\delta^{\prime}))$.
Then, we have the output of DistanceQuery (Algorithm 8) is $(\epsilon,\delta+\delta^{\prime})$-DP.
@@
The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem 3.12. Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
From advanced composition in Theorem 3.12, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
@@
Lemma E.2 (Accuracy of DistanceQuery, Algorithm 8).
@@
We follow the similar idea in the proof of Lemma D.5. Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma D.5) for the $i$-th coordinate. We can observe that the overall error across $d$ coordinates can be upper bounded by
@@
Finally, from Lemma A.3, we have the additive error given by
@@
E.2 High Dimension Single Data Structure

### v2

@@ D.3 Algorithm of Data Structure
Lemma D.3 (Runtime of initialization, Algorithm 2).
@@
Lemma D.4 (Runtime of query, Algorithm 2).
@@
Lemma D.5 (Privacy of query, Algorithm 2).
The output process of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
@@
Suppose that our dataset is $X\in[-R,R]^{n}$. Note that we only add noise in the pre-processing stage. There is no noise in the query stage. Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma C.3). Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension. We will then blow up the DP parameter by $\log n$ factor. Thus, using the basic composition rule (Fact B.6), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP. ∎
@@
Lemma D.6 (Accuracy of query, Algorithm 2).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 2.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
@@
We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$. From Fact 2.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
@@
Using Lemma B.3, we can have
@@ D.4 Disjoint Intervals
Lemma D.7 (Sum of disjoint interval queries).
@@
Then, we have Algorithm 2 is $(\epsilon,\delta)$-DP and outputs $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ with the error upper bounded by
@@
From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP. Then, from Fact B.7 and the disjoint intervals in Algorithm 6, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2 for interval $S_{j}$. Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$. From the proof of Lemma D.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
@@
Using Lemma B.3, we can have error bounded by
@@
Lemma D.8 (Weighted sum of disjoint interval queries, formal version of Lemma C.2).
@@
Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
@@
The DP proof is the same as in the proof of Lemma D.7.
Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma D.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$. From the proof of Lemma D.7, we know the variance of $z_{j}$ is bounded by
@@
Using Lemma B.3, we can have error bounded by
@@
Appendix E Weighted $\ell_{p}^{p}$ Distance
In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea. In Section E.1, we show how to solve one dimensional weighted problem. In Section E.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
@@ Appendix E Weighted $\ell_{p}^{p}$ Distance
E.1 One Dimensional Weighted Distance
@@
Lemma E.1 (Weighted distance one dimension).
@@
E.2 High Dimensional Weighted Distance
@@
Lemma E.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma C.4).
@@
where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma E.1.
Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
In this section, we generalize the algorithms in BLM+ [10] to weighted distance. Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e. $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$. In Section F.1, we analyze the runtime of our algorithm. In Section F.2, we analyze the DP and accuracy of our algorithm. In Section F.3, we give the theorem for our DPTreeDistance data structure.
@@ Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
@@
3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 2
@@
7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma E.1
@@
13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 2, Lemma C.3
@@
1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma F.2, Lemma F.4, Lemma F.5
3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 5
@@ F.1 Runtime Analysis
Lemma F.1 (Runtime of initialization, Algorithm 5).
@@
Lemma F.2 (Runtime of DistanceQuery, Algorithm 6).
@@
Remark F.3.
In Line 8 and 13 of Algorithm 6, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma E.1, i.e. $|k-j|R/n$. This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
F.2 Privacy and Accuracy Analysis
@@
Lemma F.4 (Privacy of DistanceQuery, Algorithm 6).
The output process of DistanceQuery (Algorithm 6) is $(\epsilon,\delta)$-DP.
@@
From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP output. We observe that intervals in Algorithm 6 are disjoint. Then, following the same logic in the proof of Lemma D.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP. ∎
@@
Lemma F.5 (Accuracy of DistanceQuery, Algorithm 6).
@@
From Lemma D.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
@@
Then from the proof of Lemma D.8, we can know that the variance is given by
@@
since the sensitivity $\Delta=2R_{w}$ from Lemma C.3.
Using Lemma B.3, we can have additive error bounded by
@@
F.3 One Dimension Single Data Structure
@@
Theorem F.6 (DPTreeDistance data structure, formal version of Theorem C.6).
There is a data structure DPTreeDistance (Algorithm 5,6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
@@
The proofs follow from combining Lemma F.1 (running time of initialization), Lemma F.2 (running time of query), Lemma F.4 (DP of query), and Lemma F.5 (error of query) together. ∎
Appendix G High-Dimensional Weighted $\ell_{1}$ Query
In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from BLM+ [10]. In Section G.1, we give the analysis of Algorithm 7. In Section G.2, we give the theorem of our DPTreeHighDim data structure.
Algorithm 5,6 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function. We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
G.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
@@
1:datastrucutre DPTreeHighDim $\triangleright$ Theorem G.3
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 5
@@
7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma E.2
@@
11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 5
@@
14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma G.1, Lemma G.2
@@
17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 6
@@
Lemma G.1 (Privacy of DistanceQuery, Algorithm 7).
@@
  • Let $\epsilon=O(\log(1/\delta^{\prime}))$.
Then, we have the output process of DistanceQuery (Algorithm 7) is $(\epsilon,\delta+\delta^{\prime})$-DP.
@@
The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem B.8. Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
From advanced composition in Theorem B.8, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
@@
Lemma G.2 (Accuracy of DistanceQuery, Algorithm 7).
@@
We follow the similar idea in the proof of Lemma F.5. Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma F.5) for the $i$-th coordinate. We can observe that the overall error across $d$ coordinates can be upper bounded by
@@
Finally, from Lemma B.3, we have the additive error given by
@@
G.2 High Dimension Single Data Structure